A bank password consists of two letters of the English alphabet followed by two digits. How many different passwords are there? 


1
Expert's answer
2021-10-10T18:23:54-0400

Taking into account that English Alphabet consists of 26 letters ant there are 10 digits, we conclude by Combinatorial Product Rule that the number of different passwords is equal to "26\cdot 26\cdot 10\cdot 10=67600."